OpenStudy (dannyrod2000):

convert the function 20 Degrees Celsius to Degrees Fahrenheit. F(c)=9/5c+32 20c=______F

OpenStudy (mathstudent55):

You mean convert the temperature 20 C to F?

OpenStudy (dannyrod2000):

yes

OpenStudy (mathstudent55):

Use the conversion function you have, and replace c with 20. What do you get for F(c) ?

OpenStudy (mathstudent55):

\(F(c) = \dfrac{9}{5}c + 32\) Let c = 20, so you get: \(F(c) = \dfrac{9}{5} \times 20 + 32\) Evaluate the right side.

OpenStudy (dannyrod2000):

so 9/5 which is 1.8 times 20+32?

OpenStudy (mathstudent55):

correct

OpenStudy (dannyrod2000):

just to know how do you put the fraction bar because i use a slash /

OpenStudy (dannyrod2000):

the answer would be 52 right?

OpenStudy (mathstudent55):

Use the equation editor or LaTeX commands.

OpenStudy (mathstudent55):

No. What do you get for 1.8 * 20 ?

OpenStudy (dannyrod2000):

36?

OpenStudy (mathstudent55):

Correct. Now add that to 32. What do you get?

OpenStudy (dannyrod2000):

o i messed up i added 20 and 32

OpenStudy (mathstudent55):

Right. You need to add 36 and 32 instead.

OpenStudy (dannyrod2000):

68

OpenStudy (mathstudent55):

Correct.

OpenStudy (dannyrod2000):

thanks

OpenStudy (mathstudent55):

wlcm